You were given ferrous sulfate by your doctor to help prevent iron deficiency anemia. The chemical formula of ferrous sulfate is FeSO4. How many moles are there in 115.45 g? (atomic mass: Fe = 55.85 AMU; S= 32.01 AMU; O=16.00 AMU)


1
Expert's answer
2022-01-20T05:47:01-0500

Molar mass of "FeSO_4=55.85+32.01+4(16)=151.86"


Moles = "\frac{115.45}{151.56}=0.76moles"